OpenStudy (anonymous):

At a sand and gravel plant, sand is falling off a conveyor and onto a conical pile at a rate of 10 cubic ft/ min. The diameter of the base of the cone is approximately 3x the altitude. At what rate is the height of the pile changing when the pile is 15 ft high?

OpenStudy (anonymous):

we can do this if you tell me what the formula for the volume of a cone is

OpenStudy (anonymous):

ok i found it it is \[V=\frac{1}{3}\pi r^2h\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and "diameter of base is 3 times h" means \[D=3h\] or \[r=\frac{3h}{2}\]right? so we can write everything in terms of h which is what we want

OpenStudy (anonymous):

\[V=\frac{\pi}{3}(\frac{3h}{2})^2h\]

OpenStudy (anonymous):

\[V=\frac{\pi}{3}\frac{9h^3}{4}\] \[V=\frac{3\pi h^3}{4}\] correct my algebra if it is wrong, because that is the hard part and the rest is easy

OpenStudy (anonymous):

look ok?

OpenStudy (anonymous):

its correct so far

OpenStudy (anonymous):

will we b solving thsi w/ derivatives? bc taht is how i need to solve this

OpenStudy (anonymous):

then we are almost done. take the derivative and get \[V'=\frac{9\pi h^2}{4}h'\]

OpenStudy (anonymous):

it that exponent a 3 or a 2 for base h

OpenStudy (anonymous):

you are told that \[V'=10\] it was 3, when we take the derivative it is 2

OpenStudy (anonymous):

is v' teh same as (dv/dt)' ?

OpenStudy (anonymous):

you want \[h'\] so solve \[10=\frac{9\pi h^2}{4}h'\] for \[h'\] get \[h'=\frac{40}{9\pi h^2}\]

OpenStudy (anonymous):

yes maybe you wrote in class \[\frac{dV}{dt}=\frac{dV}{dh}\times \frac{dh}{dt}\] which i wrote in shorthand as \[V'=\frac{9\pi h^2}{4}h'\]

OpenStudy (anonymous):

ok thnks 4 cleaering taht up

OpenStudy (anonymous):

now that you have \[h'=\frac{40}{9\pi h^2}\] you can plug in your value for h to get the answer

OpenStudy (anonymous):

so i substitute 15 right? i got 40/2025 pi or 8/ 405 pi

OpenStudy (anonymous):

i guess that should be right

OpenStudy (anonymous):

ok thnk u soooo much =)

OpenStudy (anonymous):

yw